OpenStudy (anonymous):

Karen is a saleswoman. She is paid a base salary of $1700 a month plus a 2% commission on the sales she makes. What was her sales amount last month if she earned a total of $2099.68 ?

OpenStudy (anonymous):

amount earned = $2099.68 base salary = $1700 difference = $399.68 let the sales amount be x 2% of x = 399.68 2/100* x =399.68 x = 399.68*100/2 = 19984